OpenStudy (anonymous):

I need more math help! What is the value of the expression below when x = -2? 5 + 3(x - 7) - (x + 1) What is 5 • 8 • (-7) What is the value of the expression 35 ÷ 7 + 8 • 3 -15? And I have another question but I have to write it in equation form.

OpenStudy (anonymous):

What is the equation for a line that passes through the points (-2,2) and (4,10) y= -2x + 6 y= 2x - 6 y= -2x + 2 y= 2x + 2 help me please

OpenStudy (agent0smith):

What is the value of the expression below when x = -2? 5 + 3(x - 7) - (x + 1) when x=-2 means replace every x with -2

OpenStudy (the_fizicx99):

5 + 3(x - 7) - (x + 1) 5 +3(-2) - 3(-2) -2 +1

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

What about the other ones

OpenStudy (the_fizicx99):

8*(-7) + and a - = - So what ever you got from 8 * (-7) multiply it by 5

OpenStudy (anonymous):

So it is -280

OpenStudy (the_fizicx99):

Yes

OpenStudy (anonymous):

Thanks! What is the value of the expression 35 ÷ 7 + 8 • 3 -15?

OpenStudy (the_fizicx99):

35 / 7 the what you got from that multiply 8*3 and add those two together and subtract -15

OpenStudy (anonymous):

OK I divided 35 ÷ 7 then I ?? I am confused

OpenStudy (the_fizicx99):

Multiply 8 *3 then add what you got from 35/7

OpenStudy (anonymous):

24 + 5 = 29

OpenStudy (anonymous):

Then 29 - 15?

OpenStudy (the_fizicx99):

Yesh

OpenStudy (anonymous):

Thank you!

OpenStudy (the_fizicx99):

:3